OpenStudy (anonymous):

a line perpendicular to a line whose slope is -0.75 has a slope of 1.3. True of false?

OpenStudy (anonymous):

m=-0.75=-3/4 perpendicular = 4/3

myininaya (myininaya):

oops true

OpenStudy (anonymous):

4/3 is not exactly 1.3 , but could be a proximate... :)

myininaya (myininaya):

well it depends is 1.3 is the 3 repeating?

myininaya (myininaya):

if the 3 is repeating then true if not then false

myininaya (myininaya):

you can tell if the 3 is repeating if they have a bar over it